# How many measurements are enough? Bayesian recovery in inverse problems with general distributions

### v1

@@ 4.1 Generative DNNs
This result shows that $\mathcal{P}$ has low complexity, since $\log(\mathrm{Cov}_{\eta,\delta}(\mathcal{P}))$ scales log-linearly in $k$. Notice that $L$ only appears logarithmically in (4.1). While it is not the main focus of this work, we note that Lipschitz constants of DNNs have been studied quite extensively fazlyab2019efficient ; szegedy2013intriguing ; virmaux2018lipschitz . Moreover, it is also possible to design and train DNNs with small Lipschitz constants miyato2018spectral .
This result, combined with Theorem 3.5 shows that accurate and stable Bayesian recovery with such a prior with sample complexity that is near-optimal in the latent dimension $k$, i.e. $\mathcal{O}(k\log(k))$.
  • Remark 4.2 (Quadratic bottleneck)
    In Theorem 3.8, the measurement condition (3.5) also depends on the coherence. This quantity has been considered in berk2022coherence for the case of ReLU DNNs. In particular, if a ReLU DNN $G:\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ has random weights drawn from a standard normal distribution, then its coherence $\mu(U;D)$ scales like $\mathcal{O}(k)$ up to log factors (berk2022coherence, , Thm. 3). Combining this with Theorem 3.8 and Proposition 4.1, we see that the overall sample complexity for Bayesian recovery scales like $\mathcal{O}(k^{2}\log(k))$ in this case. This is worse than the subgaussian case, where there is no coherence factor. The quadratic bottleneck that occurs when considering subsampled orthogonal matrices also arises in the non-Bayesian setting, as was discussed in detail in berk2022coherence . Its removal is an open problem (see §5). Note also that the coherence is not guaranteed to be small for general (in particular, trained) DNNs. However, berk2022coherence also discuss strategies for training generative models to have small coherence.
@@ 4.2 Distributions of sparse vectors
Let $s\in\{1,\ldots,n\}$. We define a distribution $\mathcal{P}=\mathcal{P}_{s}$ of $s$-sparse vectors in $\mathbb{R}^{n}$ as follows. To draw $x\sim\mathcal{P}$, we first choose a support set $S\subseteq\{1,\ldots,n\}$, $|S|=s$, uniformly at random amongst all possible (${n\choose s}$ such subsets. We then define $x_{i}=0$, $i\notin S$, and for each $i\in S$ we draw $x_{i}$ randomly and independently from $\mathcal{N}(0,1)$. Note that there are other ways to define distributions of sparse vectors, which can be analyzed similarly. However, for brevity we only consider the above setup.
@@
  • Remark 4.4 (Quadratic bottleneck)
    Once more we witness a quadratic bottleneck. In the non-Bayesian setting, one can show stable and accurate recovery of sparse vectors from $\mathcal{O}(s)$ measurements, up to log terms (see, e.g., (adcock2021compressive, , Cor. 13.15)). However, this requires specialized theoretical techniques that heavily leverage the structure of the set of sparse vectors. In the setting of this paper, the bottleneck arises from the generality of the approach considered in this work: specifically, the fact our main results hold for arbitrary probability distributions $\mathcal{P}$.
@@ 5 Limitations and future work
Second, as noted in Remarks 4.1 and 4.2 there is a quadratic bottleneck when considering subsampled orthogonal transforms. In the non-Bayesian case, this can be overcome in the case of (structured) sparse models using more technical arguments adcock2024unified . We believe similar ideas could also be exploited in the Bayesian setting. On a related note, both bora2017compressed ; berk2022coherence also consider ReLU generative models in the non-Bayesian setting, and derive measurement conditions that do not involve the Lipschitz constant of the DNN. It is unclear whether these results can be extended to the Bayesian setting.
Third, our main result involves the density shift bound (Definition 2.5). In particular, the noise distribution should have a density. This is a rather unpleasant technical assumption. It would be interesting to see if it could be removed through a refined analysis.
@@
Finally, our results in this paper are theory, and strive to study the sample complexity of Bayesian recovery. We do not address the practical problem of sampling from the posterior. This is a key computational challenge in Bayesian inverse problems arridge2019solving . However, efficient techniques for doing this are emerging. See song2022solving ; jalal2021robust and references therein.
@@
BA was acknowledges support from the Natural Sciences and Engineering Research Council of Canada of Canada (NSERC) through grant RGPIN-2021-611675.
@@
Hence $x\in B_{\eta}(x_{i})$. Since $z$ was arbitrary, we deduce that $G(B_{\eta/L}(x_{i}))\subseteq B_{\eta}(z_{i})$. It follows that $B_{\eta/L}(x_{i})\subseteq G^{-1}(B_{\eta}(z_{i}))$. Now suppose that $\gamma\left[\bigcup^{k}_{i=1}B_{\eta/L}(x_{i})\right]\geq 1-\delta$. Then, by definition of the pushforward measure
@@
Using standard properties of covering numbers (see, e.g., [4, Lem. 13.22], we get
@@
Appendix B Concentration inequalities and the proofs of Theorems 3.5 and 3.8
We now aim to prove Theorems 3.5 and 3.8. To do this, we first derive concentration inequalities for subsgaussian random matrices and subsampled orthogonal transforms.
@@ B.2 Subgaussian concentration inequalities
Let $\mathcal{A}$ be a distribution of subgaussian random matrices with parameters $\beta,\kappa>0$ (Definition 3.4). Then the lower, upper and absolute concentration bounds for $\mathcal{A}$ (Definition 2.3) can be taken as
@@
We now use [36, Lem. 9.8]. Note that this result only considers a bound of the form $\mathbb{P}(|{\|Ax\|}^{2}-{\|x\|}^{2}|\geq s{\|x\|}^{2})$ for $s\in(0,1)$. But the proof straightforwardly extends to $s>0$. ∎
@@ B.3 Concentration inequalities for randomly-subsampled orthogonal transforms
Let $D\subseteq\mathbb{R}^{n}$ and $\mathcal{A}$ be a distribution of randomly-subsampled orthoognal transforms based on a matrix $U$ (Definition 3.6). Then the lower and upper concentration bounds for $\mathcal{A}$ (Definition 2.3) can be taken as
@@
for $s>0$. The result uses Bernstein’s inequality for bounded random variables (see, e.g., [4, Thm. 12.18]). Let $x\in D$. By definition of $A$ and the fact that $U$ is orthogonal, we can write
@@
Let $p=\mathbb{P}\left[{\|x^{*}-\hat{x}\|}\geq 34(\eta+\sigma)\right]$. We use Theorem 3.1 with $c=32$, $c^{\prime}=2$ and $t=2$. Let $\varepsilon^{\prime}=\varepsilon/\delta^{1/p}$. Then Theorem 3.1 gives
@@
$+2D_{\mathsf{shift}}(2\varepsilon^{\prime},2\sigma;\mathcal{E})\mathrm{e}^{k}\left[C_{\mathsf{low}}\left(1/2;\mathcal{A},\mathbb{R}^{n}\right)+C_{\mathsf{upp}}\left(2;\mathcal{A},\mathbb{R}^{n}\right)+2D_{\mathsf{upp}}(2\sigma;\mathcal{E})\right],$
@@
$D_{\mathsf{shift}}(2\varepsilon^{\prime},2\sigma;\mathcal{E})=\exp\left(\frac{m(4\sigma+2\varepsilon^{\prime})}{2\sigma^{2}}2\varepsilon^{\prime}\right)\lesssim 1,$
where we used the facts that $m\geq 1$ and $\sigma\geq\varepsilon/\delta^{1/p}$. We deduce that
$p\lesssim\delta+\exp(k-c(\beta,\kappa)m)$
for a possibly different constant $c(\beta,\kappa)>0$. The condition (3.4) on $m$ and(3.1) now give that $p\lesssim\delta$, as required. ∎
@@
Let $p=\mathbb{P}\left[{\|x^{*}-\hat{x}\|}\geq 34(\eta+\sigma)\right]$. In this case, the forwards operator $A$ satisfies ${\|A\|}\leq\sqrt{n/m}$ (Lemma B.4). Hence we may apply Theorem 1.1 with $\theta=\sqrt{n/m}$ and $d=2$ to obtain
@@
To prove these theorems, we first require some additional background on couplings, along with several lemmas. We then require a series of technical lemmas.
@@ C.1 Background on couplings
$\gamma(E)=\sum_{s\in S,s\in\pi_{2}(E)}\gamma(E_{s}\times\{s\})+\gamma(\hat{E}),$
where $\hat{E}=E\backslash\bigcup_{s\in S,s\in\pi_{2}(E)}(E_{s}\times\{s\})$. It suffices to show that $\gamma(\hat{E})=0$. Since $F\subseteq\pi^{-1}_{2}(\pi_{2}(F))$ for any set $F$, we have
@@ C.2.1 Separation lemma
The lemma considers a scenario where two random variables are drawn for a mixture of $k$ probability distributions. The second random variable is conditioned on the draw of the first. It then considers the probability that the two random variables are drawn from different distributions in the mixture, bounding this in terms of their Total Variation (TV) distance. It generalizes [47, Lem. 3.1].
@@
Let $\mathcal{H}_{1},\ldots,\mathcal{H}_{k}$ be Borel probability measures and consider the mixture $\mathcal{H}=\sum^{k}_{i=1}a_{i}\mathcal{H}_{i}$. Let $y^{*}\sim\mathcal{H}$ and $\hat{y}\sim\mathcal{H}(\cdot|y^{*})$. Then
$\mathbb{P}[y^{*}\sim\mathcal{H}_{i},\hat{y}\sim\mathcal{H}_{j}]\leq 1-\mathrm{TV}(\mathcal{H}_{i},\mathcal{H}_{j}).$
To clarify, in this lemma and elsewhere we use the notation $y^{*}\sim\mathcal{H}_{i}$ (and similar) to mean the event that $y^{*}$ is drawn from the $i$th distribution $\mathcal{H}_{i}$. We also write $\mathcal{H}(\cdot|y^{*})$ for the posterior distribution.
@@
Note that we may assume without loss of generality that $\hat{y}$ is conditionally independent of the event $y^{*}\sim\mathcal{H}_{i}$. In this case, $\hat{y}$ follows the mixture distribution $\sum^{k}_{i=1}\mathbb{P}(y^{*}\sim\mathcal{H}_{i}|y^{*})\mathcal{H}_{i}(\cdot)$, where $\mathbb{P}(y^{*}\sim\mathcal{H}_{i}|y^{*})$ are the posterior weights. Note that if the $\mathcal{H}_{i}$ have densities $h_{i}$ with respect to some measure, these weights are given by
@@
$p:$ $=\mathbb{P}[y^{*}\sim\mathcal{H}_{i},\hat{y}\sim\mathcal{H}_{j}]$
$=\mathbb{P}[y^{*}\sim\mathcal{H}_{i}]\mathbb{P}[y^{*}\sim\mathcal{H}_{i}|\hat{y}\sim\mathcal{H}_{j}]$
$=\mathbb{P}[y^{*}\sim\mathcal{H}_{i}]\mathbb{E}[\mathbb{P}(\hat{y}\sim\mathcal{H}_{j}|y^{*})|y^{*}\sim\mathcal{H}_{i}]$
$=a_{i}\mathbb{E}[\mathbb{P}(\hat{y}\sim\mathcal{H}_{j}|y^{*})|y^{*}\sim\mathcal{H}_{i}].$
@@
$p=a_{i}\mathbb{E}[\mathbb{P}(\hat{y}\sim\mathcal{H}_{j}|y^{*})|y^{*}\sim\mathcal{H}_{i}]=a_{i}\int\mathbb{P}(\hat{y}\sim\mathcal{H}_{j}|y^{*})\,\mathrm{d}\mathcal{H}_{i}(y^{*}).$
@@
$p=a_{i}\int\mathbb{P}(\hat{y}\sim\mathcal{H}_{j}|y^{*})h_{i}(y^{*})\,\mathrm{d}\mathcal{H}(y^{*}).$
By definition, we have $\mathbb{P}(\hat{y}\sim\mathcal{H}_{j}|y^{*})=\mathbb{P}(y^{*}\sim\mathcal{H}_{j}|y^{*})$ and using (C.2), we deduce that
$p=\int\frac{a_{i}a_{j}h_{i}(y^{)}h_{j}(y^{*})}{\sum^{k}_{l=1}a_{l}h_{l}(y^{*})}\,\mathrm{d}\mathcal{H}(y^{*})$

### v2

@@ 4.1 Generative DNNs
This result shows that $\mathcal{P}$ has low complexity, since $\log(\mathrm{Cov}_{\eta,\delta}(\mathcal{P}))$ scales log-linearly in $k$. Combined with Theorem 3.5, it shows that accurate and stable Bayesian recovery with such a prior with a sample complexity that is near-optimal in the latent dimension $k$, i.e., $\mathcal{O}(k\log(k))$.
Notice that $L$ only appears logarithmically in (4.1). While it is not the main focus of this work, we note that Lipschitz constants of DNNs have been studied quite extensively fazlyab2019efficient ; szegedy2013intriguing ; virmaux2018lipschitz . Moreover, it is also possible to design and train DNNs with small Lipschitz constants miyato2018spectral .
In Theorem 3.9, the measurement condition (3.5) also depends on the coherence. This quantity has been considered in berk2022coherence for the case of ReLU DNNs. In particular, if a ReLU DNN $G:\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$ has random weights drawn from a standard normal distribution, then its coherence $\mu(U;D)$ scales like $\mathcal{O}(k)$ up to log factors (berk2022coherence, , Thm. 3). Combining this with Theorem 3.9 and Proposition 4.1, we see that the overall sample complexity for Bayesian recovery scales like $\mathcal{O}(k^{2}\log(k))$ in this case. This is worse than the subgaussian case, where there is no coherence factor and the sample complexity, as noted above, is $\mathcal{O}(k\log(k))$. Such a quadratic bottleneck also arises in the non-Bayesian setting berk2022coherence . Its removal is an open problem (see §5). Note that the coherence is also not guaranteed to be small for general (in particular, trained) DNNs. However, berk2022coherence also discuss strategies for training generative models to have small coherence. Numerically, they show that generative models with smaller coherence achieve better recovery from the same number of measurements than those with larger coherence.
@@ 4.2 Distributions of sparse vectors
Let $s\in\{1,\ldots,n\}$. We define a distribution $\mathcal{P}=\mathcal{P}_{s}$ of $s$-sparse vectors in $\mathbb{R}^{n}$ as follows. To draw $x\sim\mathcal{P}$, we first choose a support set $S\subseteq\{1,\ldots,n\}$, $|S|=s$, uniformly at random amongst all possible $n\choose s$} such subsets. We then define $x_{i}=0$, $i\notin S$, and for each $i\in S$ we draw $x_{i}$ randomly and independently from $\mathcal{N}(0,1)$. Note that there are other ways to define distributions of sparse vectors, which can be analyzed similarly. However, for brevity we only consider the above setup.
@@
Once more we witness a quadratic bottleneck. In the non-Bayesian setting, one can show stable and accurate recovery of sparse vectors from $\mathcal{O}(s)$ measurements, up to log terms (see, e.g., (adcock2021compressive, , Cor. 13.15)). However, this requires specialized theoretical techniques that heavily leverage the structure of the set of sparse vectors. In the setting of this paper, the bottleneck arises from the generality of the approach considered in this work: specifically, the fact that our main results hold for arbitrary probability distributions $\mathcal{P}$.
@@ 5 Limitations and future work
Second, as noted in Remarks 4.1 and 4.2 there is a quadratic bottleneck when considering subsampled orthogonal transforms. In the non-Bayesian case, this can be overcome in the case of (structured) sparse models using more technical arguments adcock2024unified . We believe similar ideas could also be exploited in the Bayesian setting. On a related note, both bora2017compressed ; berk2022coherence consider ReLU generative models in the non-Bayesian setting, and derive measurement conditions that do not involve the Lipschitz constant $L$ of the DNN. It is unclear whether analogous results can be established in the Bayesian setting.
Third, our main result involves the density shift bound (Definition 2.5). In particular, the noise distribution should have a density. This rather unpleasant technical assumption stems from Lemma C.6, which is a key step in proving the main result, Theorem 3.1. This lemma allows one to replace the ‘real’ distribution $\mathcal{R}$ in the probability term $p$ in Theorem 3.1 by the approximate distribution $\mathcal{P}$. This is done in order to align the prior and the posterior, which is necessary for the subsequent steps of the proof of Theorem 3.1. It would be interesting to see if this assumption on the noise could be removed through a refined analysis.
@@
Finally, our results in this paper are theoretical, and strive to study the sample complexity of Bayesian recovery. We do not address the practical problem of sampling from the posterior. This is a key computational challenge in Bayesian inverse problems arridge2019solving . However, efficient techniques for doing this are emerging. See, e.g., song2022solving ; kawar2021snips ; jalal2021robust and references therein. We believe an advantage of our results is their independence from the choice of posterior sampling algorithm, whose analysis can therefore be performed separately. This is an interesting problem we intend to examine in the future. In future work we also intend present numerical experiments for various practical settings that further support the theory developed in this paper.
@@
BA acknowledges the support of the Natural Sciences and Engineering Research Council of Canada of Canada (NSERC) through grant RGPIN-2021-611675. NH acknowledges support from an NSERC Canada Graduate Scholarship. Both authors would like to thank Paul Tupper and Weiran Sun for helpful comments and feedback.
@@
Hence $z\in B_{\eta}(z_{i})$. Since $z$ was arbitrary, we deduce that $G(B_{\eta/L}(x_{i}))\subseteq B_{\eta}(z_{i})$. It follows that $B_{\eta/L}(x_{i})\subseteq G^{-1}(B_{\eta}(z_{i}))$. Now suppose that $\gamma\left[\bigcup^{k}_{i=1}B_{\eta/L}(x_{i})\right]\geq 1-\delta$. Then, by definition of the pushforward measure
@@
Using standard properties of covering numbers (see, e.g., [5, Lem. 13.22], we get
@@
Appendix B Concentration inequalities and the proofs of Theorems 3.5 and 3.9
We now aim to prove Theorems 3.5 and 3.9. To do this, we first derive concentration inequalities for subsgaussian random matrices and subsampled orthogonal transforms.
@@ B.2 Subgaussian concentration inequalities
Let $\mathcal{A}$ be a distribution of subgaussian random matrices with parameters $\beta,\kappa>0$ (Definition 3.4). Then the lower ad upper concentration bounds for $\mathcal{A}$ (Definition 2.3) can be taken as
@@
We now use [39, Lem. 9.8]. Note that this result only considers a bound of the form $\mathbb{P}(|{\|Ax\|}^{2}-{\|x\|}^{2}|\geq s{\|x\|}^{2})$ for $s\in(0,1)$. But the proof straightforwardly extends to $s>0$. ∎
@@ B.3 Concentration inequalities for randomly-subsampled orthogonal transforms
Let $D\subseteq\mathbb{R}^{n}$ and $\mathcal{A}$ be a distribution of randomly-subsampled orthognal transforms based on a matrix $U$ (Definition 3.7). Then the lower and upper concentration bounds for $\mathcal{A}$ (Definition 2.3) can be taken as
@@
for $s>0$. The result uses Bernstein’s inequality for bounded random variables (see, e.g., [5, Thm. 12.18]). Let $x\in D$. By definition of $A$ and the fact that $U$ is orthogonal, we can write
@@
Let $p=\mathbb{P}\left[{\|x^{*}-\hat{x}\|}\geq 34(\eta+\sigma)\right]$. We use Theorem 3.1 with $c=32$, $c^{\prime}=2$, $t=2$ and $\varepsilon$ replaced by $\varepsilon/d$, where $d\geq 1$ is a constant that will be chosen later. Let $\varepsilon^{\prime}=\varepsilon/(d\delta^{1/p})$. Then Theorem 3.1 gives
@@
$+2D_{\mathsf{shift}}(2\varepsilon^{\prime},2\sigma;\mathcal{E})\mathrm{e}^{k}\left[C_{\mathsf{low}}\left(1/2;\mathcal{A},\mathbb{R}^{n}\right)+C_{\mathsf{upp}}\left(2;\mathcal{A},\mathbb{R}^{n}\right)+2D_{\mathsf{upp}}(2\sigma;\mathcal{E})\right].$
@@
$D_{\mathsf{shift}}(2\varepsilon^{\prime},2\sigma;\mathcal{E})=\exp\left(\frac{m(4\sigma+2\varepsilon^{\prime})}{2\sigma^{2}}2\varepsilon^{\prime}\right)\leq\exp\left(\frac{6m}{d}\right),$
where we used the facts that $\sigma\geq\varepsilon/\delta^{1/p}=d\epsilon^{\prime}$ and $d\geq 1$. We deduce that
$p\lesssim\delta+\exp(k+6m/d-c(\beta,\kappa)m)$
for a possibly different constant $c(\beta,\kappa)>0$. We now choose $d=d(\beta,\kappa)=12/c(\beta,\kappa)$. Up to another possible change in $c(\beta,\kappa)$, the condition (3.4) on $m$ and (3.1) now give that $p\lesssim\delta$, as required. ∎
@@
Let $p=\mathbb{P}\left[{\|x^{*}-\hat{x}\|}\geq 34(\eta+\sigma)\right]$. In this case, the forwards operator AA satisfies ${\|A\|}\leq\sqrt{n/m}$ (Lemma B.4). Hence we may apply Theorem 1.1 with $\theta=\sqrt{n/m}$ and $d=2$ to obtain
@@
We finally consider the proofs of the two general results, Theorems 1.1 and 3.1. Our main effort will be in establishing the latter, from which the former will follow after a short argument.
To prove Theorem 3.1, we first require some additional background on couplings, along with several lemmas. This is given in §C.1. We then establish a series of key technical lemmas, presented in §C.2, which are used in the main proof. Having shown these, the proof of Theorem 3.1 proceeds in §C.3 via a series of step. We now briefly describe these steps, and by doing so explain how the sets $D_{1},D_{2}$ defined in (3.2)-(3.3) arise.
  1. (i)
    First, using Lemma C.7, we decompose $\mathcal{P},\mathcal{R}$ into distributions $\mathcal{P}^{\prime},\mathcal{R}^{\prime}$ that are supported in balls of a given radius, plus remainder terms. The distributions $\mathcal{P}^{\prime},\mathcal{R}^{\prime}$ are close (in $W_{\infty}$) to a discrete distribution $\mathcal{Q}$ supported at the centres of the balls that give the approximate cover satisfying (3.1).
  2. (ii)
    Next, we replace $x^{*}\sim\mathcal{R}$ in the definition of the probability $p$ in Theorem 3.1 by $z^{*}\sim\mathcal{P}^{\prime}$. The is done to align the prior with the posterior $\mathcal{P}(\cdot|y,A)$, which is needed later in the proof. We do this using Lemma C.6. Here we have to consider the action of $A$ on vectors of the form xzx-z, where $x\in\mathrm{supp}(\mathcal{R}^{\prime})$ and $z\in\mathrm{supp}(\mathcal{P}^{\prime})$. After determining the supports of $\mathcal{R}^{\prime},\mathcal{P}^{\prime}$, we see that $x-z\in D_{1}$, where $D_{1}$ is the set defined in (3.2).
  3. (iii)
    We now decompose $\mathcal{P}^{\prime}$ into a mixture over the balls mentioned in (i). After a series of arguments, we reduce the task to that of considering the probability that the conditional distribution is drawn from one ball when the prior is drawn from another. Lemmas C.4 and C.5 handle this. They involve estimating the action of $A$ on vectors xzx-z, where $z$ is the centre of one of the balls and $x$ comes from another ball. Since the balls are supported in $\mathrm{supp}(\mathcal{P})$ we have $x\in\mathrm{supp}(\mathcal{P})$, and since the centres come from the approximate covering number bound (3.1), we have that $z\in\mathrm{supp}(\mathcal{P})$ if $\mathcal{P}$ attains the minimum and $z\in\mathrm{supp}(\mathcal{R})$ otherwise. Hence $x-z\in D_{2}$, with $D_{2}$ as in (3.3).
@@ C.1 Background on couplings
[…]
@@ C.2.1 Separation lemma
The lemma considers a scenario where two random variables are drawn for a mixture of $k$ probability distributions. The second random variable is conditioned on the draw of the first. It then considers the probability that the two random variables are drawn from different distributions in the mixture, bounding this in terms of their Total Variation (TV) distance. It generalizes [49, Lem. 3.1].
@@
Let $\mathcal{H}_{1},\ldots,\mathcal{H}_{k}$ be Borel probability measures and consider the mixture $\mathcal{H}=\sum^{k}_{i=1}a_{i}\mathcal{H}_{i}$. Let $y^{*}\sim\mathcal{H}$ and $\hat{y}\sim\sum_{i=1}^{k}\mathbb{P}(y^{*}\sim\mathcal{H}_{i}|y^{*})\mathcal{H}_{i}(\cdot|y^{*})$ where $\mathbb{P}(y^{*}\sim\mathcal{H}_{i}|y^{*})$ are the posterior weights. Then
$\mathbb{P}[y^{*}\sim\mathcal{H}_{i},\hat{y}\sim\mathcal{H}_{j}(\cdot|y^{*})]\leq 1-\mathrm{TV}(\mathcal{H}_{i},\mathcal{H}_{j}).$
To clarify, in this lemma and elsewhere we use the notation $y^{*}\sim\mathcal{H}_{i}$ (and similar) to mean the event that $y^{*}$ is drawn from the $i$th distribution $\mathcal{H}_{i}$.
@@
Note that if the $\mathcal{H}_{i}$ have densities $h_{i}$ with respect to some measure, these weights are given by
@@
$p:=\mathbb{P}[y^{*}\sim\mathcal{H}_{i},\hat{y}\sim\mathcal{H}_{j}(\cdot|y^{*})]$ $=\mathbb{P}[y^{*}\sim\mathcal{H}_{i}]\mathbb{P}[\hat{y}\sim\mathcal{H}_{j}(\cdot|y^{*})|y^{*}\sim\mathcal{H}_{j}]$
$=\mathbb{P}[y^{*}\sim\mathcal{H}_{i}]\mathbb{E}[\mathbb{P}(\hat{y}\sim\mathcal{H}_{j}(\cdot|y^{*})|y^{*}\sim\mathcal{H}_{i}]$
$=a_{i}\mathbb{E}[\mathbb{P}(\hat{y}\sim\mathcal{H}_{j}(\cdot|y^{*}))|y^{*}\sim\mathcal{H}_{i}].$
@@
$p=a_{i}\mathbb{E}[\mathbb{P}(\hat{y}\sim\mathcal{H}_{j}(\cdot|y^{*}))|y^{*}\sim\mathcal{H}_{i}]=a_{i}\int\mathbb{P}(\hat{y}\sim\mathcal{H}_{j}(\cdot|y^{*}))\,\mathrm{d}\mathcal{H}_{i}(y^{*}).$
@@
$p=a_{i}\int\mathbb{P}(\hat{y}\sim\mathcal{H}_{j}(\cdot|y^{*}))h_{i}(y^{*})\,\mathrm{d}\mathcal{H}(y^{*}).$
By definition, we have $\mathbb{P}(\hat{y}\sim\mathcal{H}_{j}(\cdot|y^{*}))=\mathbb{P}(y^{*}\sim\mathcal{H}_{j}|y^{*})$ and using (C.2), we deduce that
$p=\int\frac{a_{i}a_{j}h_{i}(y^{*})h_{j}(y^{*})}{\sum^{k}_{l=1}a_{l}h_{l}(y^{*})}\,\mathrm{d}\mathcal{H}(y^{*})$